# Ecotransformer: Attention without Multiplication

### v1

@@ 1 Introduction
Despite the remarkable success of Transformer models across a broad range of AI applications, the standard dot-product attention mechanism remains a major computational bottleneck. The core cost arises from the $QK^{T}$ matrix multiplication, which scales as $O(N^{2}D_{k})$, where $N$ is the sequence length and $D_{k}$ is the hidden dimension. For example, in GPT-3, with a sequence length of 2048 and a hidden size of 12,288, each token generation involves attention computations across 96 heads and 96 layers, resulting in billions of floating-point operations per forward pass. This computational intensity translates into substantial energy demands. In early 2023, ChatGPT was estimated to consume approximately 564 megawatt-hours per day, equivalent to the daily electricity usage of about 18,000 U.S. households (Luo and Sun,, 2024). In more extreme cases, the annual energy consumption of Google’s AI services has been estimated at 29.3 terawatt-hours, comparable to the electricity consumption of an entire country such as Ireland (De Vries,, 2023). This motivates the development of more efficient alternatives to dot-product attention that can achieve comparable expressiveness with reduced computational overhead. In this manuscript, we introduce a novel method of attention score calculation that entirely eliminates the need for matrix multiplication, departing from the traditional dot-product formulation at the core of standard Transformer architectures.
@@ 2 Distance-Based Attention Score
and $\mathbf{Q}_{i},\mathbf{K}_{j},\mathbf{V}_{j}$ are the rows of the query, key and value matrices, respectively. The dot product $<\mathbf{Q}_{i},\mathbf{K}_{j}>$ can be reformulated in terms of $L_{2}$ distance:
$<\mathbf{Q}_{i},\mathbf{K}_{j}>=\frac{1}{2}\{||\mathbf{Q}_{i}||_{2}^{2}+||\mathbf{K}_{j}||_{2}^{2}|-||\mathbf{Q}_{i}-\mathbf{K}_{j}||_{2}^{2}\}.$
@@
$\alpha_{ij}=\frac{e^{-\frac{1}{2/\sqrt{D_{k}}}||\mathbf{Q}_{i}-\mathbf{K}_{j}||_{2}^{2}}}{\sum_{j^{\prime}}e^{-\frac{1}{2/\sqrt{D_{k}}}||\mathbf{Q}_{i}-\mathbf{K}_{j^{\prime}}||_{2}^{2}}}.$
@@ 3 The Convolution Mechanism
[…]
@@ 4 Computational Complexity
In comparison, the $ij$th entry of $QK^{T}$ is
$<\mathbf{Q}_{i},\mathbf{K}_{j}>=\sum_{m=1}^{D_{k}}Q_{im}*K_{jm},$
@@ 6.1 NLP applications
By evaluating across these varied tasks, we compare the performance of the proposed $L_{1}$ attention mechanism with that of the standard scaled dot-product attention, focusing on accuracy under different types of linguistic and logical reasoning demands. All models are trained under identical settings—batch size, embedding size, sequence length, number of heads, number of layers, learning rate, and number of epochs—to ensure a fair comparison. Table 1 shows that the $L_{1}$-based method with $\lambda=1$ achieves performance very close to that of the dot-product baseline. The observed accuracy difference ranges from 0.0018-0.0018- 0.0018 to 0.0110-0.0110- 0.0110, indicating minimal degradation. Moreover, when the optimal $\lambda$ is selected via grid search, the $L_{1}$ method outperforms the dot-product method on three out of the four benchmarks. Although the optimal $\lambda$ varies across tasks, setting $\lambda=3$ yields robust performance across all datasets. In this configuration, the accuracy differences relative to the dot-product method are +0.0190+0.0190+ 0.0190, 0.0102-0.0102- 0.0102, 0.0026-0.0026- 0.0026, and +0.0034+0.0034+ 0.0034 on SciQ, StoryCloze, HellaSwag, and BoolQ, respectively. These results suggest that a properly tuned $L_{1}$ attention mechanism can achieve performance comparable to, and in some cases better than, the scaled dot-product attention across a range of NLP tasks. This highlights its potential as an efficient alternative without compromising accuracy.
@@ 6.2 Biological and Vision Applications
To evaluate the generalizability and efficiency of the proposed transformer architecture, we conduct comprehensive comparative studies across three biologically diverse datasets: The Cancer Genome Atlas (TCGA), METABRIC, and a TCR–epitope classification dataset derived from VDJdb. TCGA and METABRIC are large-scale cancer genomics cohorts that provide transcriptomic profiles and clinical outcome data, widely used for survival analysis and biomarker discovery in oncology. TCGA encompasses multi-omics data from multiple cancer types, while METABRIC focuses specifically on breast cancer and includes high-quality molecular subtyping and long-term survival annotations. In contrast, the TCR–epitope dataset captures adaptive immune recognition by pairing T-cell receptor (TCR) sequences with known antigenic epitopes, forming a sequence-level classification task with applications in immunotherapy and vaccine design. Input features include log-normalized TPM gene expression for genomics datasets and amino acid sequences for the TCR task. Additionally, we evaluate our architecture on a standard computer vision benchmark, CIFAR-10, which comprises 60,000 color images (32$\times$32 pixels) from 10 object categories, split into 50,000 training and 10,000 test samples. Across all datasets, we compare our $L_{1}$-based transformer to the standard scaled dot-product attention transformer as a baseline.
We frame each task as a multi-class classification problem: predicting cancer subtypes for TCGA and METABRIC, epitope classes from TCR sequences, and object classes for CIFAR-10. All models are trained using cross-entropy loss and evaluated on consistent data splits. To ensure fair comparison, we use identical settings across experiments, including batch size, learning rate, and number of training epochs. For evaluation, we adopt standard metrics for multi-class classification, including overall accuracy, macro-averaged precision, recall, F1 score, and area under the ROC curve (AUROC). As shown in Table 2, when the $L_{1}$-based transformer is tuned with an appropriate $\lambda$ value, it consistently outperforms the dot-product-based transformer across all datasets. The performance gains are observed across all metrics: improvements range from 0.010.010.01 to 0.250.250.25 in precision, 0.010.010.01 to 0.060.060.06 in recall, 0.010.010.01 to 0.080.080.08 in F1 score, 0.010.010.01 to 0.050.050.05 in accuracy, and from 0.02-0.02- 0.02 to 0.050.050.05 in AUROC.

### v2

@@ 1 Introduction
Despite the remarkable success of Transformer models across a broad range of AI applications, the standard dot-product attention mechanism remains a major computational bottleneck. The core cost arises from the $QK^{T}$ matrix multiplication, which scales as $O(N^{2}D_{k})$, where NN is the sequence length and $D_{k}$ is the hidden dimension. For example, in GPT-3, with a sequence length of 2048 and a hidden size of 12,288, each token generation involves attention computations across 96 heads and 96 layers, resulting in billions of floating-point operations per forward pass. This computational intensity translates into substantial energy demands. In early 2023, ChatGPT was estimated to consume approximately 564 megawatt-hours per day, equivalent to the daily electricity usage of about 18,000 U.S. households (Luo and Sun,, 2024). In more extreme cases, the annual energy consumption of Google’s AI services has been estimated at 29.3 terawatt-hours, comparable to the electricity consumption of an entire country such as Ireland (De Vries,, 2023). This motivates the development of more efficient alternatives to dot-product attention that can achieve comparable expressiveness with reduced computational overhead. In this manuscript, we introduce a novel method of attention score calculation that entirely eliminates the need for matrix multiplication, departing from the traditional dot-product formulation at the core of standard Transformer architectures.
@@ 2 Distance-Based Attention Score
and $\mathbf{Q}_{i},\mathbf{K}_{j},\mathbf{V}_{j}$ are the rows of the query, key and value matrices, respectively. The dot product $\mathbf{Q}_{i},\mathbf{K}_{j}>$ can be reformulated in terms of $L_{2}$ distance:
$\mathbf{Q}_{i},\mathbf{K}_{j}>=\frac{1}{2}\{||\mathbf{Q}_{i}||_{2}^{2}+||\mathbf{K}_{j}||_{2}^{2}|-||\mathbf{Q}_{i}-\mathbf{K}_{j}||_{2}^{2}\}.$
@@
[…]
@@ 3 The Convolution Mechanism
We examine how the $L_{2}$ and $L_{1}$ distance metrics affect attention weights differently. Consider a single embedding dimension where the query and key differ by an absolute value of $|d|$. Under the $L_{2}$-based attention mechanism, the attention weight is reduced by a multiplicative factor of $e^{-d^{2}/(2\sqrt{D_{k}})}$, whereas under the $L_{1}$-based mechanism, it is reduced by $e^{-\lambda|d|/\sqrt{D_{k}}}$. In the exponential scale, the $L_{1}$ metric results in a linear decay with respect to $|d|$, while the $L_{2}$ metric induces a quadratic decay. Figure 1 illustrates the decay functions $k(d)$ for different attention mechanisms, plotted against the absolute difference $|d|$. Compared to $L_{1}$ with $\lambda=1$, the $L_{2}$ function penalizes larger distances more severely, causing attention to focus on tokens that are very close in representation space, i.e., highly dependent tokens. In contrast, the $L_{1}$-based function distributes attention more broadly, allowing influence from tokens that are moderately or even weakly related. As the parameter $\lambda$ increases, the $L_{1}$ decay curve approaches the shape of the $L_{2}$ curve. In particular, when $\lambda=1/2D_{k}^{1/4}$, the $L_{1}$ and $L_{2}$ curves intersect at the inflection point of the $L_{2}$ curve. In this case, the two attention functions exhibit similar behaviors over a range of distances.
@@ 4 Computational Complexity
In comparison, the ijijth entry of $QK^{T}$ is
$\mathbf{Q}_{i},\mathbf{K}_{j}>=\sum_{m=1}^{D_{k}}Q_{im}*K_{jm},$
@@ 6.1 NLP applications
By evaluating across these varied tasks, we compare the performance of the proposed $L_{1}$ attention mechanism with that of the standard scaled dot-product attention, focusing on accuracy under different types of linguistic and logical reasoning demands. All models are trained under identical settings—batch size, embedding size, sequence length, number of heads, number of layers, learning rate, and number of epochs—to ensure a fair comparison. Table 1 shows that the $L_{1}$-based method with $\lambda=1$ achieves performance very close to that of the dot-product baseline. The observed accuracy difference ranges from 0.0018-0.0018 to 0.0110-0.0110, indicating minimal degradation. Moreover, when the optimal $\lambda$ is selected via grid search, the $L_{1}$ method outperforms the dot-product method on three out of the four benchmarks. Although the optimal $\lambda$ varies across tasks, setting $\lambda=3$ yields robust performance across all datasets. In this configuration, the accuracy differences relative to the dot-product method are +0.0190+0.0190, 0.0102-0.0102, 0.0026-0.0026, and +0.0034+0.0034 on SciQ, StoryCloze, HellaSwag, and BoolQ, respectively. These results suggest that a properly tuned $L_{1}$ attention mechanism can achieve performance comparable to, and in some cases better than, the scaled dot-product attention across a range of NLP tasks. This highlights its potential as an efficient alternative without compromising accuracy.
@@ 6.2 Biological and Vision Applications
To evaluate the generalizability and efficiency of the proposed transformer architecture, we conduct comprehensive comparative studies across three biologically diverse datasets: The Cancer Genome Atlas (TCGA), METABRIC, and a TCR–epitope classification dataset derived from VDJdb. TCGA and METABRIC are large-scale cancer genomics cohorts that provide transcriptomic profiles and clinical outcome data, widely used for survival analysis and biomarker discovery in oncology. TCGA encompasses multi-omics data from multiple cancer types, while METABRIC focuses specifically on breast cancer and includes high-quality molecular subtyping and long-term survival annotations. In contrast, the TCR–epitope dataset captures adaptive immune recognition by pairing T-cell receptor (TCR) sequences with known antigenic epitopes, forming a sequence-level classification task with applications in immunotherapy and vaccine design. Input features include log-normalized TPM gene expression for genomics datasets and amino acid sequences for the TCR task. Additionally, we evaluate our architecture on a standard computer vision benchmark, CIFAR-10, which comprises 60,000 color images ($\times32$ pixels) from 10 object categories, split into 50,000 training and 10,000 test samples. Across all datasets, we compare our $L_{1}$-based transformer to the standard scaled dot-product attention transformer as a baseline.
We frame each task as a multi-class classification problem: predicting cancer subtypes for TCGA and METABRIC, epitope classes from TCR sequences, and object classes for CIFAR-10. All models are trained using cross-entropy loss and evaluated on consistent data splits. To ensure fair comparison, we use identical settings across experiments, including batch size, learning rate, and number of training epochs. For evaluation, we adopt standard metrics for multi-class classification, including overall accuracy, macro-averaged precision, recall, F1 score, and area under the ROC curve (AUROC). As shown in Table 2, when the $L_{1}$-based transformer is tuned with an appropriate $\lambda$ value, it consistently outperforms the dot-product-based transformer across all datasets. The performance gains are observed across all metrics: improvements range from $0.01$ to $0.25$ in precision, $0.01$ to $0.06$ in recall, $0.01$ to $0.08$ in F1 score, $0.01$ to $0.05$ in accuracy, and from 0.02-0.02 to $0.05$ in AUROC.